Solve for each of the following equations for the given variable:
y= 4x-5 for x
w= (4/7)h +3 for h
P=a+b+c for b

Answer by checkley79(3341)

y= 4x-5 for x
4x=y+5
x=(y+5)/4 ans.
w= (4/7)h +3 for h
w=4h/7+3
w-3=4h/7
4h=7(w-3)
4h=7(W-3)/4 ans.
P=a+b+c for b
b=p-a-c ans.